At year-end 2016, total assets for Arrington Inc. were $1.7 million and accounts payable were $360,000. Sales, which in 2016 were $2.9 million, are expected to increase by 30% in 2017. Total assets and accounts payable are proportional to sales, and that relationship will be maintained; that is, they will grow at the same rate as sales. Arrington typically uses no current liabilities other than accounts payable. Common stock amounted to $485,000 in 2016, and retained earnings were $340,000. Arrington plans to sell new common stock in the amount of $155,000. The firm's profit margin on sales is 5%; 65% of earnings will be retained. What were Arrington's total liabilities in 2016? Write out your answer completely. For example, 25 million should be entered as 25,000,000. Round your answer to the nearest cent. $ How much new long-term debt financing will be needed in 2017? Write out your answer completely. For example, 25 million should be entered as 25,000,000. Do not round your intermediate calculations. Round your answer to the nearest cent. (Hint: AFN - New stock = New long-term debt.) $
Answer #1

A)Total equity = common stock +Retained earning

              = 485000+340000

              = 825000

Total asset = Total liabilities +Total equity

1,700,000 = Total liabilities + 825,000

Total liabilities = 1,700,000 - 825,000

                    = 875,000

B)

2017
Sales [2,900,000 (1+.30)] 3770000
Profit margin [sales *%of margin] [3770000*5%]

188500

Addition to retained earning in 2017 [188500*65%] 122525
2017
Total asset [1,700,000(1+.30)] 2,210,000
Accounts payable [360000(1+.30)] 468,000
2017
common stock
Beginning 485000
Addition during the year 155000 640000
Retained earning
Beginning 340000
Addition during the year 122525 462525
Total equity at end of 2017 1,102,525

Total asset = Total liabilities +Total equity

2,210,000 = Total liabilities + 1,102,525

Total liabilities in 2017 = 2,210,000 -1,102,525

                   = 1,107,475

Total liabilities =current liabilities + Long term debt

1107475 = 468000 +long term debt

long term debt = 1107475 -468000

               = 639475

Long term debt at end of 2017 =Beginning long term debt + addition long term debt financing

639475 = 515000+addition long term debt financing

addition long term debt financing = 639475 -515000

                            = 124475

**Beginning long term debt = Total liabilities at beginning -current liabilities

                 = 875000 --360000

                    = 515000